Calculate the energy, in kilojoules, of one mole of photons of blue laser light of wavelength 435 nm.
Answer #1

Given:

lambda = 4.35*10^-7 m

Find energy of 1 photon first

Given:

lambda = 4.35*10^-7 m

use:

E = h*c/lambda

=(6.626*10^-34 J.s)*(3.0*10^8 m/s)/(4.35*10^-7 m)

= 4.57*10^-19 J

This is energy of 1 photon

Energy of 1 mol = energy of 1 photon * Avogadro's number

= 4.57*10^-19*6.022*10^23 J/mol

= 2.752*10^5 J/mol

= 2.752*10^2 KJ/mol

Answer: 275 KJ